OpenStudy (anonymous):

Renee can sew a 25-square-foot quilt in 16 hours. Genera can sew the same size quilt in 12 hours. How long would it take them to sew the quilt working together?

OpenStudy (anonymous):

options are 1. 6 hours 2. 6 6/7 hours 3. 8 6/7 hours 4. 28 hours

OpenStudy (***[isuru]***):

hi, first u have to find the rate at which renee saw the quilt rate = 25 / 16 square - foot quilt per hour then do the same for Genera rate = 25/12 square - foot quilt per hour If they work together the the will able to saw an area of ( 25/16 + 25/12 ) quilt per hour now the total area to saw is 25 .. and we know the rate at both of them will saw... So to find the duration .. we have to divide the area by the rate at which they saw.. \[25\div(\frac{ 25 }{ 16 } +\frac{ 25 }{ 12 })\] \[\frac{ 25(12\times16) }{ 25(12 + 16) }\]\[\frac{ 192 }{ 28 } = 6 \frac{ 6 }{ 7 }\] Hope this will help ya!!!

OpenStudy (anonymous):

thank you!

OpenStudy (***[isuru]***):

u r welcome!!